What is the formula for number of images formed by two plane mirrors when they are inclened at some particular angle

Grade:12

6 Answers

Himanshu
8 years ago
when two mirrors are arranged at an angle theta then we evaluate the quantity 360/theta and find the following cases1. 360/theta=n=even then no of images =n-1
2.n is odd integer and object placed at bisector then number of images = n-1
Gaurav Kumar
7 years ago
1.if 360/ theta is even ,no is 360/theta-1 ....2.if 360/theta is odd nd object is at angle bisector then also no is 360/theta-1 .....3.if 360/theta is odd nd object is not at angle bisector then no is 360/theta.
Sristy
6 years ago
CASE 2:​If1.   360 / \Theta = n = Odd2.  And object placed at bisector of the angle between the mirrorsthen only number of images = n-1then  number of images =  n - 1When two mirrors are arranged at an angle theta then we evaluate the quantity360 / \Theta = nand find the following casesCASE 1: IF1.  360 / \Theta = n = even
CASE 3 :
 
1.   360 / \Theta = n = Odd2.    And object is NOT placed at bisector of the angle between the mirrorsthen only number of images = n
 
A
6 years ago
 the formula for number of images formed by two plane mirrors when they are inclened at some particular angle
     1.       360/theta-1 if there is an odd angle
     2.       360/theta only if there is an even angle
Abhishek Kundu
6 years ago
1:When 360/theta is even then no. of image = 360/theta-1...2:when 360/theta is odd then no. of image = 360/theta -1(for object on the angel bisector)......3:360/theta=odd then no. of image =360/theta(for object anywhere else).
Mani Prakash Tiwary
5 years ago
360/theta =fractions , number of image = integers part of 360/theta , 360/theta odd and object at asymmetrical place image = 360/theta ; 360/theta =odd and object at symmetrical place image=360/theta-1;360/theta= even  number of image= 360/theta -1